OpenStudy (anonymous):

Simplify. -9+(-5)x+3+(-2)x

OpenStudy (anonymous):

Whats the : -9+3= ? Answer this please :)

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

-6 - 7x?

OpenStudy (anonymous):

No ! Now answer this : -5x-2x = ?

OpenStudy (anonymous):

no its not ause thats not one of the choices

OpenStudy (anonymous):

3 – 16x –12 – 7x –6 + 3x –6 – 7x

OpenStudy (anonymous):

its one of those^^^

OpenStudy (anonymous):

Look we can just sum the same things for example we cant sum the 7x and -6 . OK ? Now we get in xs and numbers : -5x-2x = -7x -6 OK ? So answer is -7x-6 Got it ?:)

OpenStudy (anonymous):

ok then im gonna go wih the last one

OpenStudy (anonymous):

wat ???

OpenStudy (anonymous):

wich one is it then???